OpenStudy (anonymous):

Use the product, quotient, and power rules of logarithms to rewrite the expression as a single logarithm. Assume that all variables represent positive real numbers. 4log x + 2log y

OpenStudy (anonymous):

First you need to move the 4 and the 2 and show them as exponents of x and y. Can you show me that?

OpenStudy (anonymous):

so x^4 and y^2

hartnn (hartnn):

yes, correct, now use the property that \(\Large \log A+\log B = \log AB\) so, whats, \(\Large \log x^4+\log y^2 = ... ?\)

OpenStudy (anonymous):

logx^4y^2?

OpenStudy (anonymous):

Exactly.

hartnn (hartnn):

yes, correct :)